# Slicing Wasserstein over Wasserstein via Functional Optimal Transport

### v1

@@ B.1 Metric Properties
. To prove the positive definiteness of our double-sliced metric for empirical meta-measures, we utilize an extension of the ‘Cramer–Wold’ theorem by Cuesta-Albertos et al. (2007) about the set of projections required to separate measures on $\mathbb{R}^{d}$. The statement is based on the so-called Carleman condition. A measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ fulfills this condition if all moments
@@
This condition is fulfilled for compactly supported measures (Schmüdgen et al., 2017, Ch. 14) and, in particular, empirical measures. We also refer to (Heppes, 1956; Tanguy et al., 2024) for similar results targeted at empirical measures.
@@ B.2 Relationships between Metric
Now, we state the last result. Note that our proof is inspired by a proof in (Piening & Beinert, 2025b).
@@
pointwisely. Since $\mathcal{X}$ is compact, $(\bm{\mu}_{n_{m}})_{m\in\mathbb{N}}$ can be chosen such that $\mu_{n_{m},k}$ converges to some $\tilde{\mu}_{k}~\in~\mathcal{P}_{2}(\mathbb{R}^{d})$. Moreover, since $\bm{\mu}_{n}$ and $\bm{\mu}$ are both empirical meta-measures with $N$ support points, we can assume without loss of generality that $\bm{\gamma}^{n,*}_{\theta,k,\ell}$ is a permutation matrix, i.e., $\bm{\gamma}^{n,*}_{\theta,k,\ell}~\in~\{0,\frac{1}{N}\}^{N\times N}$ with only one positive value per row or column (Peyré & Cuturi, 2019, Prop. 2.1).
@@ C.3 Sliced Functional Optimal Transport on $L_{2}([0,1])$
Here, we want to showcase the properties of the $\xi$-based SW distance on $\mathcal{P}_{2}(L_{2}([0,1]))$ by looking at increasingly finer function discretizations for different kernel parameters $\sigma$. We consider two empirical measure pair in $\mathcal{P}_{2}(L_{2}([0,1]))$ defined via
@@
C.4 Extension of Section 5.3 on Point Cloud Comparison
@@
Figure 10: Averaged WoW and DSW estimates between sets of point clouds for $10$ to $100$ shapes and projection number $S=100$, $S=1,000$, $S=5,000$.
@@
Figure 12: Averaged WoW and DSW estimates between sets of point clouds with $100$ to $1000$ points per shape and projection number $S=1000$, $S=10,000$, $S=50,000$.
C.5 Extension of Section 5.4 on Image Comparison via Patches
C.5.1 Representing Images via Patches
@@
C.5.2 Additional Experimental Details
In the experiment from Section 5.4, we compare distributions over synthetic texture images. We visualize samples from our random Perlin texture model (Perlin, 1985) in Figure 14. Note that our images with varying lacunarity (14(a)) are all generated with the following Perlin parameters: persistence of $1$, scale of $100$, 66 octaves. The generation model will be released with the code. For our images with varying persistence (14(b)), we use different Perlin parameters: lacunarity of $2.5$, scale of $100$, 55 octaves. Note that while the resulting images in Figure 14(a) and Figure 14(b) look rather similar, the ones from Figure 14(a) display a higher blur and less high-frequency artifacts.
@@ C.5.2 Additional Experimental Details
Figure 14: Samples from our Perlin texture noise for varying lacunarity (14(a)) and ‘persistence’ (14(b)).
@@
Figure 16: Comparing synthetic texture image batches via Euclidean Wasserstein, our sliced patch-based distance, patch-based WoW, and the KID for varying ‘lacunarity’ (6(a)) and ‘persistence’ (16(b)).

### v2

@@ B.1 Metric Properties
To prove the positive definiteness of our double-sliced metric for empirical meta-measures, we utilize an extension of the ‘Cramer–Wold’ theorem by Cuesta-Albertos et al. (2007) about the set of projections required to separate measures on $\mathbb{R}^{d}$. The statement is based on the so-called Carleman condition. A measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ fulfills this condition if all moments
@@
This condition is fulfilled for compactly supported measures (Schmüdgen and others, 2017, Ch. 14) and, in particular, empirical measures. We also refer to (Heppes, 1956; Tanguy et al., 2024) for similar results targeted at empirical measures.
@@ B.2 Relationships between Metric
Now, we state the last result. Note that our proof is inspired by a proof in (Piening and Beinert, 2025b).
@@
pointwisely. Since $\mathcal{X}$ is compact, $(\bm{\mu}_{n_{m}})_{m\in\mathbb{N}}$ can be chosen such that $\mu_{n_{m},k}$ converges to some $\tilde{\mu}_{k}~\in~\mathcal{P}_{2}(\mathbb{R}^{d})$. Moreover, since $\bm{\mu}_{n}$ and $\bm{\mu}$ are both empirical meta-measures with $N$ support points, we can assume without loss of generality that $\bm{\gamma}^{n,*}_{\theta,k,\ell}$ is a permutation matrix, i.e., $\bm{\gamma}^{n,*}_{\theta,k,\ell}~\in~\{0,\frac{1}{N}\}^{N\times N}$ with only one positive value per row or column (Peyré and Cuturi, 2019, Prop. 2.1).
@@ C.1 Implementation Details
C.2 Practical Guide on Parameter Choice
The numerical implementation depends on multiple parameters, namely the kernel parameter $\sigma$, the grid size $R$, and the projection number $S$. We conduct multiple parameter studies that can be found in Sections C.3, C.5, and C.6.
For the parameters $S$ and $R$, we note that both of them are essentially integration parameters of a Monte Carlo integration. The parameter $R$ is the size of the quantile integration grid, and $S$ is the number of Monte Carlo steps. Thus, we would generally expect better results for higher choices of $R$ and $S$. However, the runtime scales linearly with the number of Monte Carlo steps $S$ and Monte Carlo typically converges with a rate of $\mathcal{O}(S^{-1/2})$. Thus, choosing $S$ is mainly a question of balancing performance and runtime, where the marginal value of an additional Monte Carlo projection decreases for high $S$. As for the grid size $R$, we note that the quantile functions we aim to integrate are rather simple. In particular, they are monotonically increasing step functions. In line with this observation, we only found small marginal performance increases for $R>10$ in most experiments. Moreover, we note that high choices of $R$ result in sampling from highly correlated, high-dimensional Gaussians, which might become numerically unstable for some implementations.
During our experiments regarding SQW for the shape experiment in Section 5.1, we only found a small impact of $\sigma$. This is in line with Proposition 3.2 showing that variations of $\sigma$ result in an equivalent metric. While $\sigma=0$ works for our discretization in practice and essentially reduces to the slicing approach discussed in (Piening and Beinert, 2025a), it corresponds to sampling a white noise process that is not contained in $L_{2}([0,1])$. Therefore, $\sigma=0$ is not covered by our theory. The limit $\sigma\to\infty$ leads to an increasing correlation and to sampling constant test functions. However, as constant functions are not dense in $L_{2}([0,1])$, the conditions of Theorem 3.1 would not be met anymore. Hence, SQW might lose its metric properties. Thus, we generally recommend choosing $\sigma$ as a value from $[0.001,0.1]$.
@@ C.3 Sliced Functional Optimal Transport on $L_{2}([0,1])$
Both of our introduced sliced distances, SQW and DSW, rely on the $\xi$-based SW distance on $\mathcal{P}_{2}(L_{2}([0,1]))$ and its numerical implementation. We study the impact of our parameter choices on this distance by looking at increasingly finer function discretizations $R$ for different kernel parameters $\sigma$. We consider two empirical measure pair in $\mathcal{P}_{2}(L_{2}([0,1]))$ defined via
@@
C.4 Extension of Section 5.1 on Shape Classification
C.4.1 MNIST-2000
MNIST-2000 used in Section 5.1 is a synthetic dataset, which we build based on the training samples of MNIST. For each of the five digits 0–4, we first compute the mean image. From each digit, we generate 20 empirical point clouds consisting each of 2000 points. For this, we sample the mean image proportionally to pixel intensity. Afterwards, we disturb the point clouds by Gaussian noise and apply a random rotation. The points clouds are equipped with Euclidean distance, which are represented by matrices. In total, this yields 20 almost isometric shapes per digit. We show one sample per class in Figure 10.
Figure 10: MNIST-2000 Samples.
C.4.2 Stability Analysis of Classification Pipeline
The shape classification experiment presented in Table 1 depends on a random initialization determined by a random seed and on the choice of the $k$-nearest-neighbor ($k$-NN) algorithm. Since each dataset includes only about 100 shapes, cross-validation offers too little statistical support for selecting $k$. For this reason, we fix $k$ a priori. We perform an additional experiment to demonstrate the impact of $k$ and the random seed.
For our main experiment, we choose $k=3$ instead of $k=1$ to mitigate the impact of small outliers. Additionally, we aim for a small $k$ due to the limited number of samples per class, e.g., 10 samples per class for the ‘2D Shapes’ dataset. To illustrate the impact of the choice of $k$ and the random initialization, we rerun our classification pipeline for the ‘2D Shapes’ and the ‘Animals’ dataset with a different initialization and report the resulting accuracies for $k=1,3,5$. The results are presented in Table 2 and are in line with those presented in Table 1 in the main text. We observe a drop in performance for $k=5$, presumably due to limited data per class.
k 1-NN Acc. (%) 3-NN Acc. (%) 5-NN Acc. (%)
SQW (Ours) $9.9\pm 0.6 99$$9.4\pm 1.3$ $98.6\pm 2.1$
TLB $100.0\pm 0.1$ $100.0\pm 0.3$ $99.9\pm 0.4$
STLB $9.8\pm 0.6 99$$9.4\pm 1.3$ $98.3\pm 2.2$
AE $9.9\pm 0.4 99$$7\pm 1.0 99.$$9.0\pm 1.9$
GW $100.0\pm 0.2$ $99.8\pm 0.6$ $98.9\pm 1.1$
(a) 2D Shapes
1-NN Acc. (%) 3-NN Acc. (%) 5-NN Acc. (%)
$9.5\pm 0.9 99$$9.1\pm 1.3$ $68.5\pm 2.1$
$\pm 0.0 100.0±0.010$$\pm 0.0$ $70.9\pm 0.0$
$9.8\pm 0.7 99$$9.3\pm 1.1$ $69.2\pm 1.7$
$98.9\pm 1.4$ $97.8\pm 1.9$ $67.6\pm 2.3$
$\pm 0.0 100.0±0.010$$\pm 0.0$ $70.9\pm 0.0$
(b) Animals
Table 2: Classification accuracies for the nearest-neighbor algorithm (‘$k$-NN Acc.’) for varying neighborhood definitions $k$ and a random seed that is different from the one used in Table 1.
C.5 Extension of Section 5.2 on Dataset Comparison
C.5.1 Relation between Classification Accuracy and Dataset Distances
Originally, the OTDD and the s-OTDD have been introduced to estimate the difficulty of transfer learning. Therefore, we extend the experiment presented in Figure 3 to consider the connection between the different dataset distances and classifier accuracy. As a simple experiment, we consider the accuracy of a 5-nearest-neighbor (5-NN) MNIST classifier as our target. In particular, for each of our 100 MNIST splits, we average the two accuracies of classifying the first split based on the second split and vice versa. Now, this averaged accuracy is our target quantity.
Again, we estimate the correlation between this target accuracy and various dataset distances. Here, we would expect a negative correlation since the dataset splits that are further away from each other should lead to worse training results. Beyond the previously reported OTDD, s-OTDD ($S=10^{5}$), and our DSW ($S=10^{5}$, 1000 outer projections, 10 inner ones), we additionally report results for the Gaussian OTDD, the s-OTDD with $S=10^{4})$, our DSW with $S=1000$ (100 outer projections, 10 inner ones). Note that the Gaussian OTDD approximates all inner pairwise Wasserstein distances via a Gaussian approximation (Alvarez-Melis and Fusi, 2020). Essentially, this distance reduces to the mixture Wasserstein distance between Gaussian mixtures (Delon and Desolneux, 2020), since we compute all distances without feature costs, i.e, on $\mathcal{P}(\mathcal{P}(\mathbb{R}^{d}))$ instead of $\mathcal{P}(\mathcal{Y}\times\mathcal{P}(\mathbb{R}^{d}))$.
Looking at the results in Figure 12, we observe the strongest correlation for the OTDD and the Gaussian OTDD with values of $\approx-0.5$. Regarding the sliced metrics, we observe a stronger correlation for our DSW ($\approx-0.45$) compared to the s-OTDD ($\approx-0.35$). As for the number of projections, both distances seem to be stable with respect to the number of random projections $S$, and we observe only small differences between $S=10^{4}$ and $S=10^{5}$.
Generally, we observe a clearly negative correlation of approximately 0.5-0.5 across metrics. While these correlations are overall weaker than in other related experiments (Alvarez-Melis and Fusi, 2020; Nguyen et al., 2025), we attribute this to the small variability of the 5-NN classifier accuracy (ranging from 88% to 94%) and the omission of feature costs. Thus, we hypothesize that the inclusion of the feature cost via hybrid hierarchical slicing (Nguyen and Ho, 2024) would further strengthen these correlations.
(a) OTDD
(b) Gaussian OTDD
(c) s-OTDD ($S=10^{4}$)
(d) s-OTDD ($S=10^{5}$)
(e) Ours ($S=10^{4}$)
(f) Ours ($S=10^{5}$)
Figure 12: Scatter plots and correlations ($\uparrow$) between the accuracy of a 5-NN MNIST classifier and the OTDD (12(b)), the Gaussian-OTDD (12(b)), the s-OTDD with $S=10^{4},10^{5}$ (12(c), 12(d)) and DSW (‘Ours’) with $S=10^{4},10^{5}$ (12(e), 12(f)). All distances are computed without a feature cost on the labels.
C.5.2 Stability Analysis of Slicing Parameters
We investigate the impact of the projection number $S$, grid size $R$ and the kernel parameter $\sigma$ by considering the correlation between DSW and OTDD on MNIST displayed in Figure 3(a). As the projection number $S$ is a product of the number of outer and inner projections, see Section C.1, we fix the number of inner projections to $10$ and only vary the number of outer projections. In particular, we simulate the DSW distance for all combinations of $S=10^{3},10^{4},10^{5}$, $R=10,100,200$, and $\sigma=0.01,0.1$.
For each simulation, we compute the Pearson and Spearman correlation coefficients and present the results in Table 3. All coefficients are approximately in the range from $0.9$ to $0.95$. We see that the number of slices $S$ is by far the most important parameter, whereas the impact of the grid size $R$ and the kernel parameter $\sigma$ is more subtle. In particular, increasing $S$ from $10^{3}$ to $10^{4}$ increases the correlation significantly (from $0.9$ to $0.94$), whereas the increase from $10^{4}$ to $10^{5}$ only has a slight impact (raising the correlation from $0.94$ to $0.95$).
$S$ $R$ $\sigma$ Pearson Spearman
1000 10 0.1 0.9182 0.9187
1000 10 0.01 0.8966 0.8908
1000 100 0.1 0.9365 0.9385
1000 100 0.01 0.9270 0.9214
1000 200 0.1 0.9265 0.9283
1000 200 0.01 0.9063 0.9133
10000 10 0.1 0.9422 0.9369
10000 10 0.01 0.9453 0.9424
10000 100 0.1 0.9400 0.9402
10000 100 0.01 0.9513 0.9474
10000 200 0.1 0.9424 0.9392
10000 200 0.01 0.9536 0.9534
100000 10 0.1 0.9463 0.9451
100000 10 0.01 0.9524 0.9510
100000 100 0.1 0.9444 0.9438
100000 100 0.01 0.9508 0.9494
100000 200 0.1 0.9442 0.9427
100000 200 0.01 0.9533 0.9515
Table 3: Correlation between our DSW and OTDD (without feature cost) for various configurations on MNIST, extending the results in Figure 3(a).
C.6 Extension of Section 5.3 on Point Cloud Comparison
@@ C.5.2 Stability Analysis of Slicing Parameters
Figure 14: Averaged WoW and DSW (‘Ours’) estimates between sets of point clouds for $10$ to $100$ shapes and projection number $S=100$, $S=1,000$, $S=5,000$.
@@
Figure 16: Averaged WoW and DSW (‘Ours’) estimates between sets of point clouds with $100$ to $1000$ points per shape and projection number $S=1000$, $S=10,000$, $S=50,000$.
C.7 Extension of Section 5.4 on Image Comparison via Patches
C.7.1 Representing Images via Patches
@@
C.7.2 The Relevance of Patch Distributions
Many advances in image processing rely on the importance of local image features (Zontak and Irani, 2011). Indeed, convolutional neural networks in computer vision repeatedly apply the same filter to a small receptive field, and this receptive field can be understood as a patch. Moreover, vision transformers decompose images into smaller patches (Dosovitskiy et al., 2021). The advantage of this approach can be motivated by the relevance of small-range dependencies within images.
Notably, a key disadvantage of the standard MSE in imaging is its vulnerability to ‘small’ image operations. A small shift of all pixels can lead to the explosion of the MSE. However, such operations have only a small effect on the patch distributions. Indeed, the same goes for the translation of an object within an image, see visualizations in (He et al., 2024). In addition, the perceived style of an image seems to be inherently linked to certain localized image features. As an example, style transfer algorithms successfully capture certain artistic aspects of painting via such features (Gatys et al., 2016). Moreover, texture images are characterized by a certain type of stationarity, where a model can generate texture images by simply matching the patch distribution of a single exemplary texture image (Houdard et al., 2023). This could explain why patch-based WoW-type methods lead to clearer discriminiation than Euclidean Wasserstein methods, cf. Figure 20.
C.7.3 Additional Experimental Details
In the experiment from Section 5.4, we compare distributions over synthetic texture images. We visualize samples from our random Perlin texture model (Perlin, 1985) in Figure 18. Note that our images with varying lacunarity (18(a)) are all generated with the following Perlin parameters: persistence of $1$, scale of $100$, 66 octaves. The generation model will be released with the code. For our images with varying persistence (18(b)), we use different Perlin parameters: lacunarity of $2.5$, scale of $100$, 55 octaves. Note that while the resulting images in Figure 18(a) and Figure 18(b) look rather similar, the ones from Figure 18(a) display a higher blur and less high-frequency artifacts.
@@ C.7.3 Additional Experimental Details
Figure 18: Samples from our Perlin texture noise for varying lacunarity (18(a)) and ‘persistence’ (18(b)).
@@
Figure 20: Comparing synthetic texture image batches via Euclidean Wasserstein, patch-based DSW (‘Ours’), patch-based WoW, and KID for varying ‘lacunarity’ (7(a)) and ‘persistence’ (20(b)).
Appendix D Use of LLMs